Solve for x using logs.
2^x=e^(x+5)
I know im supposed to use ln, but I get stuck after that.
Please help, thanks.

Answer by John10(297)

Yes, you are correct!
Take the natural log of both sides:
ln(e^(x + 5)) = ln(2^x)
x + 5 = x ln2
x + 5 = (0.69)x
x - 0.69x = -5
0.31x = -5
x = -16.13
John10:)